OpenStudy (anonymous):

Which is 10^–3 written in standard form? A. –30 B. 0.0001 C. 0.001 D. 1000

geerky42 (geerky42):

Remember. \[a^{-b} = \dfrac{1}{a^b}\] so knowing that 10^3 = 1000: 10^-3 = 1/1000 Can you convert 1/1000 to decimal?

OpenStudy (anonymous):

it is the letter d

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

sure thing only if I could get help no problem jnejls

geerky42 (geerky42):

umm it's not D.. it's C... @gianna

OpenStudy (anonymous):

OKAY sorry typo